MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2007/08/14
OBSERVER: Ohnishi Paul
CONDITION: fine
COMMENT: Some of the images are trailing in left side and rightside. This cause due to shutter close time is longer than expected. Change the setting the moa_cam3 by KOKI, it works well. Dark D2570(60sec) for GB B9676-B9700. Dark D2573(60sec) for GB B9701-B9724. Dark D2572(300sec) is OK.
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 18:57 END 06:47


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB9676-97244929617760
LMC1622-16265298700
SMC419-41912600
ALERT-000
GRB-000
TOTAL---5532727060 (07h31m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
1011101046910104691210
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
9991099999210


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
122221554111
Field13Field14Field15Field16
1100


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
110000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
18:00waitingfine
19:00observingfine
20:00observingfine
21:00observingfine
22:00observingfine
23:00observingfine
00:00observingfine
01:00observingfine
02:00observingfine
03:00observingfine
04:00observingfine
05:00observingfine
06:00observingfine
06:35waitingstop due to twilight


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day